OpenStudy (istim):

A marble rolls off a table with a velocity of 1.93m/s[horizontally]. The tabletop is 76.5cm above the floor. If air resistance is negligible, determine the velocity at impact

OpenStudy (istim):

Apparently I misinterpreted my question. So now I got a problem.

OpenStudy (turingtest):

is it 4.3

OpenStudy (istim):

I took this as a regular acceleration question, and used \[V _{2}^{2}=v _{1^{2}}+2ad\]

OpenStudy (turingtest):

good...

OpenStudy (anonymous):

how is it rolled off

OpenStudy (istim):

Yeah, its 4.33.How?

OpenStudy (istim):

No other info given. Previous questions tho.

OpenStudy (istim):

Time is 0.396, and displacement is .763m.

OpenStudy (turingtest):

solve what you have for V2 V1 is zero the number you get is the y-component of the velocity

OpenStudy (turingtest):

the x-component does not change

OpenStudy (istim):

Im suppose to break it up into x and y components?

OpenStudy (istim):

I got 5.799m/s (down)

OpenStudy (turingtest):

they are already broken up\[V_x=1.93\]because it rolls of horizontally, and that component is constant because no force acts in that direction. The vertical component is\[V_y=\sqrt{2gh}\]and the total velocities magnitude is\[|V|=\sqrt{V_x^2+V_y^2}\]

OpenStudy (turingtest):

rolls off*

OpenStudy (anonymous):

reaction force?

OpenStudy (istim):

How did you get such formulas? Both 2?

OpenStudy (turingtest):

like I said, the horizontal component is constant because there is no force acting on the object in the x-direction, so the horizontal velocity is the 1.93m/s it rolled of the table with

OpenStudy (anonymous):

can angular rotation come into play?

OpenStudy (istim):

I'm not that advd yet.

OpenStudy (anonymous):

okies

OpenStudy (turingtest):

the y-component is what your formula leads to for the up-down portion of the motion if you recognize that the initial downward velocity is zero\[\large v_y^2=v_{yi}^2+2gh=(0)^2+2gh\to v_y=\sqrt{2gh}\]

OpenStudy (turingtest):

@mth none of that stuff is necessary so why use it?

OpenStudy (anonymous):

if air resistance negligable gravity?

OpenStudy (istim):

Gravity applies.

OpenStudy (turingtest):

gravity is in my formulation

OpenStudy (anonymous):

on the rolling of the marble

OpenStudy (istim):

So I rearrange V22=v12+2ad

OpenStudy (istim):

Apparently I can't copy equations.

OpenStudy (anonymous):

i was asking

OpenStudy (turingtest):

V22=v12+2ad ^^ ^^ ^ ^ Vy2 0 g h leads to Vy2=2gh

OpenStudy (anonymous):

so potential energy does not play in part

OpenStudy (istim):

im doing potential energy later. i hope.

OpenStudy (turingtest):

V22=v12+2ad ^^ ^^ ^ ^ Vy2 0 g h *sorry got shifted around

OpenStudy (istim):

So my result is 5.79m/s. What's my next course of action.

OpenStudy (turingtest):

where did you get that number?

OpenStudy (anonymous):

sorry thats all i can do :( my math sucks i like theoretics much better

OpenStudy (istim):

(1.93^3+2(9.8)(0.763))^1/2

OpenStudy (istim):

IT's alright mth.

OpenStudy (istim):

I re-did it, and got 4.32.

OpenStudy (turingtest):

I keep trying to tell you that the equations are separate Vx=1.93 Vy=(2*9.890.763)^1/2 you have Vx plugged in where the initial speed goes, which is zero.

OpenStudy (istim):

Oh. Shoot.

OpenStudy (turingtest):

if you wanted both those things in the same equation it would have to be a vector formula, and I don't think you've done parametric stuff yet.

OpenStudy (istim):

this time i got 3.87

OpenStudy (turingtest):

\[V_x=1.93\]\[V_y=\sqrt{2gh}\]\[|V|=\sqrt{V_x^2+V_y^2}=\sqrt{1.93^2+2(9.8)(0.765)}\approx4.33\]I'm going to sleep here so I gotta wrap it up...

OpenStudy (istim):

Pythagorean Theorem. Got it.

OpenStudy (turingtest):

Vy changes independently from Vx. That is a handy thing to know. Good luck!